OpenStudy (anonymous):

all real number that are less than or equal to -8 or greather than 12

OpenStudy (tkhunny):

Any particular notation that is required? \(x \le -8\;or\;x > 12\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@baziham

OpenStudy (tkhunny):

Okay, maybe \((-\infty,-8]\cup(12,\infty)\)?

OpenStudy (anonymous):

umm id dont thank thats it

OpenStudy (anonymous):

@zaynahf

OpenStudy (anonymous):